MVC Inc. for year 2017 had $20m in revenues, and operating expenses (incl. depreciation) of $17m. With an opportunity cost of capital of 9%, and corp. tax rate of 35%, it invested $15m. How much economic value did it add during the year?

Select one:

a. $600,000

b. $500,000

c. $1.35m

d. $1.95m

e. None of the Above
Answer #1

Economic Value Added

It refers to the economic profit of a firm. In other words, Economic value addition is an excess earning of firm over the required return of the Firm.

Thus,

EconomicValueadded(EVA)- Profitaftertax-(WACC*Investedcapital)

Where,

WACC = weighted average cost of capital (opportunity cost)

We have following information -

Revenue = $ 20,000,000

Operating expenses = $ 17,000,000

Tax rate = 35%

Opportunity cost (WACC) = 9%

Profit after tax = (20,000,000 - 17,000,000)*(1-0.35)

= 3,000,000*0.65

= $ 1,950,000

Thus,

Economic value added = 1,950,000 - (15,000,000*0.09)

= 1,950,000 - 1,350,000

= $ 600,000

Therefore, option-a is correct i.e $ 600,000